OpenStudy (jellybot23):

How on earth do you solve (e^x)(e^x+1)=1

OpenStudy (mysticfury):

uhhhh

Nnesha (nnesha):

is it \[(e^x)(e^{x+1})=1\] ?

Nnesha (nnesha):

if yes then use the exponent rule when we multiply same bases we should `add` the exponents \[\large\rm x^m \times x^n=x^{m+n}\]

OpenStudy (anonymous):

$$e^{-x} = (e)(e^x)$$ $$e^{2x} = e^{-1}$$ $$x = -0.5$$

Nnesha (nnesha):

and then use the fact ln cancel out the e\[\rm \cancel{\ln e}^a = a\]

OpenStudy (jellybot23):

Yes! Oh, so the two would just combine?

OpenStudy (jellybot23):

Wait, so if you added x and x+1, then would it just become e^2x+1?

Nnesha (nnesha):

is it e^{x+1} or\[ e^x+1 \] which one is correct

OpenStudy (jellybot23):

the first one e^{x+1}

Nnesha (nnesha):

alright good and yes then it would be \[\rm e^{2x+1}=1\]

OpenStudy (jellybot23):

and then just take the ln and solve for x, correct? :)

Nnesha (nnesha):

correct

OpenStudy (jellybot23):

Okay! Thank you so so much!

Nnesha (nnesha):

np :=))

OpenStudy (jellybot23):

After solving it, I got -1/2

Nnesha (nnesha):

that's correct!! good work